Analysis and Interpretation of Profitability
Balance sheets and income statements for 3M Company follow.

Consolidated Statements of Income
Years ended December 31 ($ millions) 2010 2009 2008
Net sales $26,662 $23,123 $25,269
Operating expenses
Cost of sales 13,831 12,109 13,379
Selling, general and administrative expenses 5,479 4,907 5,245
Research, development and related expenses 1,434 1,293 1,404
Loss/(gain) from sale of business -- -- 23
Total operating expenses 20,744 18,309 20,051
Operating income 5,918 4,814 5,218
Interest expenses and income
Interest expense 201 219 215
Interest income (38) (37) (105)
Total interest expense 163 182 110
Income before income taxes 5,755 4,632 5,108
Provision for income taxes 1,592 1,388 1,588
Net income including noncontrolling interest 4,163 3,244 3,520
Less: Net income attributable to noncontrolling interest 78 51 60
Net income $ 4,085 $ 3,193 $ 3,460
Consolidated Balance Sheets
($ millions) 2010 2009
Assets
Current Assets
Cash and cash equivalents $ 3,377 $ 3,040
Marketable securities-current 1,101 744
Accounts receivable-net 3,615 3,250
Inventories
Finished goods 1,476 1,255
Work in process 950 815
Raw materials and supplies 729 569
Total inventories 3,155 2,639
Other current assets 967 1,122
Total current assets 12,215 10,795
Marketable securities-noncurrent 540 825
Investments 146 103
Property, plant and equipment 20,253 19,440
Less: Accumulated depreciation (12,974) (12,440)
Property, plant and equipment-net 7,279 7,000
Goodwill 6,820 5,832
Intangible assets-net 1,820 1,342
Prepaid pension benefits 74 78
Other assets 1,262 1,275
Total assets $ 30,156 $ 27,250
Liabilities
Current liabilities
Short-term borrowings and current portion of long-term debt $ 1,269 $ 613
Accounts payable 1,662 1,453
Accrued payroll 778 680
Accrued income taxes 358 252
Other current liabilities 2,022 1,899
Total current liabilities 6,089 4,897
Long-term debt 4,183 5,097
Pension and postretirement benefits 2,013 2,227
Other liabilities 1,854 1,727
Total liabilities 14,139 13,948
Equity
3M Company shareholders' equity: Common stock, par value $.01 per share; 9 9
Additional paid-in capital 3,468 3,153
Retained earnings 25,995 23,753
Treasury stock (10,266) (10,397)
Accumulated other comprehensive income (loss) (3,543) (3,754)
Total 3M Company shareholders' equity 15,663 12,764
Noncontrolling interest 354 538
Total equity 16,017 13,302
Total liabilities and equity $ 30,156 $ 27,250

Compute net operating profit after tax (NOPAT) for 2010. Assume that the combined federal and statutory rate is: 37.0% (Round your answer to the nearest whole number.)
2010 NOPAT = 4266 millions

Compute net operating assets (NOA) for 2010 and 2009. Treat noncurrent Investments as a nonoperating item.
2010 NOA = 16305 millions
2009 NOA =14300 millions

Compute 3M's RNOA, net operating profit margin (NOPM) and net operating asset turnover (NOAT) for 2010. (Round your answers to two decimal places. Do not round until your final answer. Do not use NOPM x NOAT to calculate RNOA.)
2010 RNOA =27.88%
2010 NOPM =16%
2010 NOAT =1.74

(a) Compute net nonoperating obligations (NNO) for 2010 and 2009.
2010 NNO =

($ millions)
2009 NNO =

($ millions)

(B) Compute return on equity (ROE) for 2010. (Round your answers to two decimal places. Do not round until your final answer.)
2010 ROE =

%

(C) What is the nonoperating return component of ROE for 2010? (Round your answers to two decimal places.)
Hint: Use your prior rounded answers to compute this answer.
2010 nonoperating return =

%
Answer #1

Part A

net nonoperating obligations (NNO) =Nonoperating Liabilities + Noncontrolling Interest – Nonoperating Assets

2010 NNO = (1269+4183) + 0 – (3377+1101+540+146) = $288

2009 NNO = (613+5097) + 0 – (3040+744+825+103) = $998

Part B

ROE = NI/Average SE = 4085/((15663+12764)/2) = 28.74%

Part C

2010 nonoperating return =ROE – RNOA =28.74%-27.88% = 0.86%